# Conditioning Diffusions Using Malliavin Calculus

### v1

@@ Abstract
In stochastic optimal control and conditional generative modelling, a central computational task is to modify a reference diffusion process to maximise a given terminal-time reward. Most existing methods require this reward to be differentiable, using gradients to steer the diffusion towards favourable outcomes. However, in many practical settings, like diffusion bridges, the reward is singular, taking an infinite value if the target is hit and zero otherwise. We introduce a novel framework, based on Malliavin calculus and path-space integration by parts, that enables the development of methods robust to such singular rewards. This allows our approach to handle a broad range of applications, including classification, diffusion bridges, and conditioning without the need for artificial observational noise. We demonstrate that our approach offers stable and reliable training, outperforming existing techniques.
@@ 1 Introduction
Simulating conditioned diffusions is a central computational task in many applications, ranging from molecular dynamics and physical chemistry (Dellago et al., 2002; Bolhuis et al., 2002; Vanden-Eijnden et al., 2010) and genetics (Wang et al., 2011) to finance, econometrics (Bladt & Sørensen, 2014; Elerian et al., 2001; Durham & Gallant, 2002), generative modelling (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) and evolutionary biology (Arnaudon et al., 2023, 2017; Baker et al., 2024).
@@
where the drift vector field $b$ is given–either from a learned generative model or from a physically or financially motivated system. Samples from (1) can be obtained by straightforward numerical simulation.
Conditioning on Observations. However, in many settings, we also have an observation
$Y=G(X_{T})$ (2)
and would like to condition (1) on the event $Y=y$ (which could fix a property of the sample, or an area of the state space). If $Y=y$ is a rare event, then simulating many paths of (1) until $y$ occurs can be very wasteful.
Instead, we augment this setup by introducing
$\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+u_{t}(X_{t})\,\mathrm{d}t+\mathrm{d}B_{t},$ (3)
so that a suitable control $u_{t}$ can redirect the trajectory in light of new data or particular constraints (e.g. conditioning on rare events or outputting a sample with a specific property).
@@
Singular Rewards. In many settings there is no gradient information available for $g$. One case where this happens is if $G$ is an indicator function or discontinuous as it would be the case for classification. But even if $G$ is smooth but deterministic, the induced likelihood $g$ is often singular if one does not add artifical noise to the observations. Even in the seemingly straightforward case of $G=\text{Id}$, the reward would be given by a Dirac delta distribution $g(x_{T};y)=\delta_{y}(x_{T})$, which is not even continuous. The case of $G=\text{Id}$ conditions diffusions to end at a specific state and is known under the term diffusion bridges. Due to $g$ being a Dirac it is in some sense the most challenging setting and will be a guiding problem for us in the development of the methodology.
@@
Outline and Contributions. Guided by 5, we develop loss functions $\mathcal{L}(u)$ whose unique minimisers $u_{t}(x;y)$ drive the controlled diffusion in 3 to form bridges of 1 for any condition $Y=y$. Towards this goal, we
  • recall the well-known (Eberle, 2015; Denker et al., 2024; Shi et al., 2024; Du et al., 2024) connection of diffusion bridges to Doob’s $h$-transform in Section 2.1, in particular the relevance of conditional score functions,
  • lift 5 to integration by parts on the space of trajectories; the role of Lebesgue measure $\mathrm{d}x$ is replaced by the law of $(X_{t})_{0\leq t\leq T}$, and the ordinary derivative $\partial_{x}$ is replaced by the Malliavin derivative (Nualart, 2006). Based on this, we derive a novel formula for conditional scores that generalises Tweedie’s formula (Efron, 2011) for denoising score matching (Vincent, 2011),
@@ 2 Theoretical Background and Main Result
0:  $\alpha:[0,1]\to\mathbb{R}^{n\times n}$, initial condition $x_{0}$, batch size $N$, current drift approximation $u^{\theta}$, time grid $\{t_{0},t_{1},\ldots t_{M}\}$. Initialize
@@
4:     Compute the Monte Carlo estimator $\mathcal{S}_{s}$ using 8 (for details see Algorithm 2) along the path $(X,B)$.
@@
7:  Summ for the full-batch loss
$\mathcal{L}^{M}(\theta)=\sum_{i=1}^{N}l_{i}(\theta).$
8:  Take a gradient step on $\mathcal{L}^{B}(\theta)$ with your favourite optimiser.
@@ 2.1 Doob’s $h$-transform and Conditional Scores
As a first step towards Theorem 2.1, we recall the fact that the diffusion bridge drift in 10 can be expressed in terms of conditional scores (Rogers & Williams, 2000, p. 83):
@@
$u_{t}^{*}(x;y)\coloneqq\sigma_{t}(x)\sigma_{t}(x)^{\top}\nabla_{x}\log p_{T\mid t}(Y=y\mid X_{t}=x),$ (11)
@@
Similar closed-form formulations are available in the context of optimal control (Nüsken & Richter, 2021, Section 2.2) and time reversals (Boffi & Vanden-Eijnden, 2024; Song et al., 2021; Ho et al., 2020). Note that for $t=T$ $\nabla_{x}\log p_{T|t}(Y=y|X_{T}=x)=\nabla_{x}\log g(x;y)$. As a consequence, if $g$ is singular, (11) becomes singular (and possibly numerically unstable) as $t\rightarrow T$. We deal with this issue in the next subsection.
@@ 2.2 Malliavin Calculus and Integration by Parts
$\begin{split}\nabla_{x}\log p_{T\mid t}&(Y=y\mid X_{t}=x)\\ &=\mathbb{E}\left[\mathcal{S}_{t}\mid X_{t}=x,Y=y\right]\end{split}$ (12)
@@
Proof sketch.
See Appendix D.1 for full details. We rely on two main ideas:
@@
Secondly, as outlined in the introduction, we elevate the integration-by-parts formula (5) to Wiener space, i.e., to the space of sample paths of Brownian motion $(B_{t})_{0\leq t\leq T}$. To this end, we introduce the Malliavin derivative $D_{t}$, which represents differentiation with respect to the infinitesimal noise increment $\mathrm{d}B_{t}$: for a functional $F$ that depends on the realisation of $(B_{t})_{0\leq t\leq T}$, the Malliavin derivative is given by
$D_{t}F=\frac{\partial F}{\partial\,\mathrm{d}B_{t}}.$ (14)
@@
$\mathbb{E}\left[\int_{0}^{T}D_{t}F\cdot g_{t}\,\mathrm{d}t\right]=\mathbb{E}\left[F\int_{0}^{T}g_{t}\cdot\mathrm{d}B_{t}\right],$ (15)
for appropriate choices of $F$ and the stochastic process $g_{t}$.
We will see in Proposition D.2 that we can express $\nabla\log p(Y=y|X_{t}=x)$ as a conditional expectation of $\nabla\log p(X_{T}=x_{T}|X_{t}=x)$. This simplifies the problem to inferring the diffusion bridge term $\nabla\log p(X_{T}=x_{T}|X_{t}=x)$, which we do now.
@@
Remark 2.4.
Note that $\nabla\log p_{t}$ contains the Lebesgue-density $p_{t}$ and therefore does not naturally carry over to infinite dimensions or manifolds. However, the expression (8) as well as (12) can also be made sense of in infinite dimensions or for diffusions on manifolds. Therefore, these are natural expressions to use for conditioning infinite as well as manifold-valued diffusions (Baker et al., 2024).
Remark 2.5.
For linear SDEs Proposition 2.3 can be simplfied to the celebrated Tweedie formula (Efron, 2011), given, e.g., by
Similarly, the denoising score matching objective for diffusion models (Song et al., 2020) can be seen as an instance of 9 for a specific choice of $\alpha$. This also allows us to derive novel regression targets for diffusion models. Lastly, similar to Remark 2.5, the obtained formulas carry over to the infinite dimensional as well as manifold-valued setting and are natural targets to generalize diffusion models to these settings (Pidstrigach et al., 2024; De Bortoli et al., 2022). For details on this, see Appendix B.
@@ 2.3 Amortised Bridges and Forward-KL
While Propositions 2.2 and 2.3 together provide an expression for the desired control vector field $u_{t}^{*}$, the left-hand side of (12) requires access to the targeted distribution of (6), conditioned on $X_{t}=x$ and $Y=y$; the construction might thus appear circular. Fortunately, taking expectations over $X_{t}$ and $Y$ in (12) not only recovers the straightforward-to-simulate reference diffusion (6) as a mixture of its bridges, but also allows us to amortise the learning procedure, inferring all the bridges simultaneously.
@@
Since $u$ is a function of $X_{t}$ and $Y$, and since the conditional expectation is the minimizer of the $L^{2}$ distance (Klenke, 2013, Corollary 8.17), the minimizer of
$\mathcal{L}^{t,y}_{\text{local}}(u):=\mathbb{E}\left[\left\|u_{t}(X_{t},Y)-\mathcal{S}_{t}\right\|^{2}\right],$ (16)
@@
arg $\text{arg min}~{}\mathcal{L}^{t}_{\text{local}}(u)=\mathbb{E}[\mathcal{S}_{t}|X_{t},Y]=\nabla_{X_{t}}\log p(Y|X_{t}).$ (17)
Here, the second equality holds because of Proposition 2.3. Finally, by Proposition 2.2, $\nabla_{X_{t}}\log p(Y|X_{t})$ is indeed equal to the control term. By exchanging the expectation with the integral in 9, we get that $u$ is equal to the optimal control $\nabla_{X_{t}}\log p(Y|X_{t})$ for each $t$.
@@ 3.1 Optimal $\alpha$ for Reduced Variance
assuming that $\frac{\alpha_{1}-\alpha_{s}}{\sqrt{1-s}}\rightarrow 0$ as $s\rightarrow 1$. The variance is minimised for the choice
@@ 3.2.2 Gaussian Approximations
$~{}~{}~{}~{}~{}\coloneqq\mathbb{E}\big{[}\|u_{s}(X_{s},X_{T})-\nabla_{X_{s}}\log p(X_{t}\mid X_{s})\|^{2}\big{]}$
is given by the diffusion bridge $u^{*}_{t}(x_{t},x_{T})$, see (11).
Thus, optimizing $\mathcal{L}^{\text{GA}}$ for each $s$ yields the diffusion bridge drift. To achieve this, one must select a $t\geq s$ for each $s$. After discretizing the time domain into $\{t_{0},t_{1}=t_{0}+\delta t,\ldots,t_{N}=t_{0}+N\delta t\}$, a suitable loss function is:
@@
$a_{t}(X_{t})$ $=\sigma_{t}(X_{t})\sigma_{t}(X_{t})^{T}$ (24b)
This provides an explicit expression for $\nabla_{X_{t}}\log p(X_{t+\delta t}\mid X_{t})$, which can be used for regression. In Heng et al. (2022), the authors applied a similar Gaussian approximation to the time-reversal of a diffusion to learn bridges of the time-reversed process.
However, this approximation is performed in the density domain $p(X_{t+\delta t}\mid X_{t})$, and its impact on the accuracy of $\nabla\log p(X_{t+\delta t}\mid X_{t})$ is unclear. By the generalized Tweedie formula (see Proposition 2.3), we obtain the explicit relation:
$\nabla\log p_{t}(X_{t+\delta t}\mid X_{t})=\mathbb{E}[\mathcal{S}_{s}\mid X_{t},X_{t+\delta t}],$ (25)
where $\mathcal{S}_{s}$ is defined in (8) with terminal time $T=t+\delta t$.The Gaussian approximation can now directly be related to a discretization scheme for $\mathcal{S}_{s}$ in (25). We prove this in the case of $\sigma=1$:
@@
Approximating $p(X_{t+\delta t}\mid X_{t})$ by a Gaussian (24a) and regressing against its score is equivalent to a choosing $\alpha^{\prime}_{s}=1_{[t,t+\delta t]}$ and approximating the stochastic integral in $\mathcal{S}_{s}$ (8) as
@@
The proofs for this section can be found in Section D.4.‡

### v2

@@ Abstract
In generative modelling and stochastic optimal control, a central computational task is to modify a reference diffusion process to maximise a given terminal-time reward. Most existing methods require this reward to be differentiable, using gradients to steer the diffusion towards favourable outcomes. However, in many practical settings, like diffusion bridges, the reward is singular, taking an infinite value if the target is hit and zero otherwise. We introduce a novel framework, based on Malliavin calculus and centred around a generalisation of the Tweedie score formula to nonlinear stochastic differential equations, that enables the development of methods robust to such singularities. This allows our approach to handle a broad range of applications, like diffusion bridges, or adding conditional controls to an already trained diffusion model. We demonstrate that our approach offers stable and reliable training, outperforming existing techniques. As a byproduct, we also introduce a novel score matching objective. Our loss functions are formulated such that they could readily be extended to manifold-valued and infinite dimensional diffusions.
@@ 1 Introduction
Simulating conditioned diffusions is a central computational task in many applications, ranging from molecular dynamics and physical chemistry (Dellago et al., 2002; Bolhuis et al., 2002; Vanden-Eijnden et al., 2010) and genetics (Wang et al., 2011) to finance, econometrics (Bladt & Sørensen, 2014; Elerian et al., 2001; Durham & Gallant, 2002), evolutionary biology (Arnaudon et al., 2023, 2017; Baker et al., 2024), and generative modelling (Dhariwal & Nichol, 2021; Ho & Salimans, 2021), including guidance for generative models (Zhang et al., 2023; Denker et al., 2024).
@@
where the drift vector field $b$ is given–either from a learned generative model or from physically or financially motivated considerations. Samples from (1) can be obtained by straightforward numerical simulation.
Guidance and Control. In many applications, it is desirable to condition the reference process (1). For instance, in generative modeling, a large pretrained model may not offer the necessary task-specific controls. One may then seek to impose such controls post hoc—for example, to generate images matching a desired edge map or human pose, or to sample proteins with properties tailored to a particular application. Similarly, when the reference process models weather dynamics, interest may lie in rare or extreme scenarios rather than typical ones.
Bayesian Inverse Problems. These objectives can be formalised within the framework of Bayesian inverse problems (Stuart, 2010). To this end, let $G$ be an observation operator and $Y$ the corresponding observation at terminal time:
$Y=G(X_{T}).$ (2)
The goal is to sample $(X_{t})_{0\leq t\leq T}$ given $Y=y$. This setting encompasses all previous examples; for instance, $G$ may extract edge or pose information from an image. Our formulation allows $G$ to be a full or partial (potentially noisy) observation of $X_{T}$. Note that $G$ may be stochastic; for example, $G(X_{T})=X_{T}+\xi$, where $\xi$ is a random variable.
Stochastic Optimal Control and Conditioning Diffusions. In order to condition on $Y$, we add a control term $u_{t}$ to the reference process (1):
$\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+u_{t}(X_{t})\,\mathrm{d}t+\mathrm{d}B_{t}.$ (3)
The optimal control will have the property that each sample $(X_{t})_{0\leq t\leq T}$ from (3) is a sample from the distribution of (1), conditioned on $Y=y$.
@@
Singular Rewards. In many settings there is no gradient information available for $g$. One case where this happens is if $G$ is an indicator function or discontinuous, as would be the case for classification. But even if $G$ is smooth but deterministic, the induced likelihood $g$ is often singular if one does not add artificial noise to the observations. Even in the seemingly straightforward case of $G=\text{Id}$, the reward would be given by a Dirac delta distribution $g(x_{T};y)=\delta_{y}(x_{T})$, which is not even continuous. The case of $G=\text{Id}$ conditions diffusions to end at a specific state and is known under the term diffusion bridges (see Figure 1). Due to $g$ being a Dirac it is in some sense the most challenging setting and will be a guiding problem for us in the development of the methodology.
@@
Outline and Contributions. Guided by 5, we develop loss functions $\mathcal{L}(u)$ whose unique minimisers $u_{t}(x;y)$ implement the conditioning of (1) via the controlled diffusion in 3, for any condition $Y=y$ at once. Towards this goal,
  • we recall the well-known (Eberle, 2015; Denker et al., 2024; Shi et al., 2024; Du et al., 2024) connection of diffusion bridges to Doob’s $h$-transform in Section 2.1, in particular the relevance of conditional score functions,
  • lift 5 to integration by parts on the space of trajectories; the role of the Lebesgue measure $\mathrm{d}x$ is replaced by the law of $(X_{t})_{0\leq t\leq T}$, and the ordinary derivative $\partial_{x}$ is replaced by the Malliavin derivative (Nualart, 2006). Based on this, we derive a novel formula for conditional scores that generalises Tweedie’s formula (Efron, 2011) for denoising score matching (Vincent, 2011),
@@ 2 Theoretical Background and Main Result
Remark 2.2 (Amortisation).
After training according to Algorithm 1, the learned control $u_{t}(x;y)$ performs conditioning of (1) for arbitrary $y$, without the need for retraining. On a related note, the constructions in (8) and (9) as well as in Algorithm 1 depend on the observation operator $G$ only implicitly through the samples $Y=G(X_{T})$ and do not access or require any particular structure; in particular, they are gradient free. On the technical technical level this corresponds to the conditional expectation being the minimiser of the $L^{2}$ loss, see (17).
@@
0:  $\alpha:[0,1]\to\mathbb{R}^{n\times n}$, initial condition $x_{0}$, batch size $N$, current drift approximation $u^{\theta}$, time grid $\{t_{0},t_{1},\ldots t_{M}\}$.
@@
4:     Compute the Monte Carlo estimator for $\mathcal{S}_{s}$ using 8 (for details see Algorithm 2) along the path $(X,B)$.
@@
7:  Sum for the full-batch loss
$\mathcal{L}^{N}(\theta)=\sum_{i=1}^{N}l_{i}(\theta).$
8:  Take a gradient step on $\mathcal{L}^{N}(\theta)$ with your favourite optimiser.
@@ 2.1 Doob’s $h$-transform and Conditional Scores
As a first step towards Theorem 2.1, we recall the fact that the optimal drift in 10 can be expressed in terms of conditional scores (Rogers & Williams, 2000, p. 83):
@@
$u_{t}^{*}(x;y)\coloneqq\sigma_{t}(x)\sigma_{t}(x)^{\top}\nabla_{x}\log p(Y=y\mid X_{t}=x),$ (11)
@@
Similar closed-form formulations are available in the context of optimal control (Nüsken & Richter, 2021, Section 2.2) and time reversals (Boffi & Vanden-Eijnden, 2024; Song et al., 2021b; Ho et al., 2020). Note that for $t=T$, the conditional score coincides with the gradient of the log-reward, $\nabla_{x}\log p(Y=y\mid X_{T}=x)=\nabla_{x}\log g(x;y)$. As a consequence, if $g$ is singular, (11) becomes singular (and possibly numerically unstable) as $t\rightarrow T$. We deal with this issue in the next subsection.
@@ 2.2 Malliavin Calculus and Integration by Parts
$\begin{split}\nabla_{x}\log p&(Y=y\mid X_{t}=x)\\ &=\mathbb{E}\left[\mathcal{S}_{t}\mid X_{t}=x,Y=y\right],\end{split}$ (12)
@@
Remark 2.5.
In the case when (6) is linear, Proposition 2.4 simplifies to the celebrated Tweedie formula (Efron, 2011), given, e.g., by
Similarly, the denoising score matching objective for diffusion models (Song et al., 2021a) can be seen as an instance of 9 for a specific choice of $\alpha$. Consequently, (12) is a generalisation to nonlinear SDEs, also allowing us to derive novel regression targets for diffusion models; see Appendix B.
Proof sketch (Proposition 2.4).
See Appendix D.1 for full details. Without loss of generality, we may consider the diffusion bridge setting (i.e., $G(X_{T})=X_{T}$), see Lemma D.2. To show (12), we rely on two main ideas:
@@
Secondly, as outlined in the introduction, we elevate the integration-by-parts formula (5) to Wiener space, i.e., to the space of sample paths of Brownian motion $(B_{t})_{0\leq t\leq T}$. To this end, we introduce the Malliavin derivative $D_{t}$, which represents differentiation with respect to the infinitesimal noise increment $\mathrm{d}B_{t}$: for a functional $g$ that depends on the realisation of $(B_{t})_{0\leq t\leq T}$, the Malliavin derivative is given by
$D_{t}g=\frac{\partial g}{\partial\,\mathrm{d}B_{t}}.$ (14)
@@
$\mathbb{E}\left[\int_{0}^{T}D_{t}g\cdot f_{t}\,\mathrm{d}t\right]=\mathbb{E}\left[g\int_{0}^{T}f_{t}\cdot\mathrm{d}B_{t}\right],$ (15)
for appropriate choices of $g$ and the stochastic process $f_{t}$.
@@ 2.3 Amortised Bridges and Forward-KL
While Propositions 2.3 and 2.4 together provide an expression for the desired control vector field $u_{t}^{*}$, the left-hand side of (12) requires access to the targeted distribution of (6), conditioned on $X_{t}=x$ and $Y=y$; the construction might thus appear circular. Fortunately, taking expectations over $X_{t}$ and $Y$ in (12) not only recovers the straightforward-to-simulate reference diffusion (6) as a mixture of its bridges, but also allows us to amortise the learning procedure, inferring all the bridges simultaneously.
@@
The minimiser of the least-squares “local-in-time” loss
$\mathcal{L}^{t}_{\text{local}}(u):=\mathbb{E}\left[\left\|u_{t}(X_{t},Y)-\mathcal{S}_{t}\right\|^{2}\right]$ (16)
@@
$u^{*}_{t}(x,y)=\mathbb{E}[\mathcal{S}_{t}\mid X_{t}=x,Y=y].$ (17)
Combining Proposition 2.3 and Proposition 2.4, this implies that the minimiser $u_{t}^{*}$ is equal to the optimal control term. By exchanging the expectation with the integral in 9, we get that $u_{t}$ is equal to the optimal control $\nabla_{X_{t}}\log p(Y\mid X_{t})$, for each $t\in[0,T]$.
@@ 3.1 Optimal $\alpha$ for Reduced Variance
assuming that $\frac{\alpha_{1}-\alpha_{s}}{\sqrt{1-s}}\rightarrow 0$ as $s\rightarrow 1$ and letting $d=x_{1}-x_{0}$. The variance is minimised for the choice
@@ 3.2.2 Gaussian Approximations
$~{}~{}~{}~{}~{}\coloneqq\mathbb{E}\big{[}\|u_{s}(X_{s},y)-\nabla_{X_{s}}\log p_{t\mid s}(X_{t}\mid X_{s})\|^{2}\big{]}$
is minimised by the diffusion bridge term $u^{*}_{s}(x_{s},y)$, see (11).
Thus, optimising $\mathcal{L}^{\text{GA}}$ for each $s$ yields the diffusion bridge drift. To achieve this, one must select a $t\geq s$ for each $s$. After discretising the time domain into $\{t_{0},t_{1}=t_{0}+\delta t,\ldots,t_{N}=t_{0}+N\delta t\}$, a suitable loss function is:
@@
$a_{t}(X_{t})$ $=\sigma_{t}(X_{t})\sigma_{t}(X_{t})^{\top}.$ (24b)
This provides an explicit expression for $\nabla_{X_{t}}\log p_{t+\delta t\mid t}(X_{t+\delta t}\mid X_{t})$, which can be used for regression. In Heng et al. (2022), the authors applied a similar Gaussian approximation to the time-reversal of a diffusion to learn bridges of the time-reversed process.
However, this approximation is performed in the density domain $p_{t+\delta t\mid t}(X_{t+\delta t}\mid X_{t})$, and its impact on the accuracy of $\nabla\log p_{t+\delta t\mid t}(X_{t+\delta t}\mid X_{t})$ is unclear. By the generalised Tweedie formula (see Proposition 2.4), we obtain the explicit relation:
$\nabla\log p_{t+\delta t\mid t}(X_{t+\delta t}\mid X_{t})=\mathbb{E}[\mathcal{S}_{s}\mid X_{t},X_{t+\delta t}],$ (25)
where $\mathcal{S}_{s}$ is defined in (8) with terminal time $T=t+\delta t$. The Gaussian approximation can now be directly related to a discretisation scheme for $\mathcal{S}_{s}$ in (25). We prove this in the case of $\sigma=1$:
@@
The proofs for this section can be found in Section D.4.